OpenStudy (anonymous):

Solve the inequality. Round your answer to the nearest tenth. X^2 + 4⋅x - 9 > 5

OpenStudy (tkhunny):

Subtract 5. Find where the new expression is zero (0). Pick the correct sides of those zeros.

OpenStudy (anonymous):

I really don't understand what you mean by that

OpenStudy (tkhunny):

Please subtract 5. You must understand this or we cannot proceed.

OpenStudy (anonymous):

okay X^2+4x-14>0 ?

OpenStudy (tkhunny):

Change to = 0 and solve for x.

OpenStudy (anonymous):

\[x ^{2}+4x-9>5\] \[x ^{2}+4x>5+9,x ^{2}+4x+4>14+4\] \[\left( x+2 \right)^{2}>18\] \[\left| x+2 \right|>3\sqrt{2}\] \[x+2<-3\sqrt{2}and x+2 > 3\sqrt{2}\] find x

OpenStudy (tkhunny):

This gives \(x = 3\sqrt{2} - 2\) and \(x = -3\sqrt{2} - 2\) as decision points. You can be very analytical about it from here or you can just try things out. What happens if we try x = -7 What happens if we try x = 0 What happens if we try x = 3

OpenStudy (anonymous):

x<-3*1.4-2 or x <-5.2-2 or x < -7.2 x >3*1.4-2 or x > 5.2-2 or x > 3.2